Wildhorse Ltd. purchased a new machine on April 4, 2014, at a cost of $152,000. The company estimated that the machine would have a residual value of $14,000. The machine is expected to be used for 9,200 working hours during its four-year life. Actual machine usage was 1,300 hours in 2014; 2,000 hours in 2015; 2,500 hours in 2016; 1,800 hours in 2017; and 1,600 hours in 2018. Wildhorse has a December 31 year end.

(a)

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is partially correct.

Calculate depreciation for the machine under each of the following methods: (Round expense per unit to 2 decimal places, e.g. 2.75 and final answers to 0 decimal places, e.g. 5,275.)

(1) Straight-line for 2014 through to 2018.

2014 expense $enter a dollar amount
2015 expense $enter a dollar amount
2016 expense $enter a dollar amount
2017 expense $enter a dollar amount
2018 expense $enter a dollar amount



(2) Diminishing-balance using double the straight-line rate for 2014 through to 2018.

2014 expense $enter a dollar amount
2015 expense $enter a dollar amount
2016 expense $enter a dollar amount
2017 expense $enter a dollar amount
2018 expense $enter a dollar amount



(3) Units-of-production for 2014 through to 2018.

2014 expense $enter a dollar amount
2015 expense $enter a dollar amount
2016 expense $enter a dollar amount
2017 expense $enter a dollar amount
2018 expense $enter a dollar amount
Answer #1

Answer-1)- The depreciation expense for 2014 = $25875.

Year 2015 depreciation expense = $34500.

Year 2016 depreciation expense = $34500.

Year 2017 depreciation expense = $34500.

Year 2018 depreciation expense = $34500.

Explanation- Straight line Method-

= Cost of asset- Salvage value of asset/No. of useful life (years)

=($152000-$14000)/4 years

=$138000/4 years

= $34500         

Year 2014 depreciation expense =( $34500*9 months)/12 months

= $25875

Year 2015 depreciation expense = $34500.

Year 2016 depreciation expense = $34500.

Year 2017 depreciation expense = $34500.

Year 2018 depreciation expense = $34500.

2)- The depreciation expense for 2014 = $57000.

The depreciation expense for 2015 = $47500.

The depreciation expense for 2016 = $23750

The depreciation expense for 2017 = $11875.

The depreciation expense for 2018 = $2125.

Explanation- Double Declining balance depreciation is calculated using the following formula=

Depreciation = Depreciation Rate * Book Value of Asset

Depreciation rate is given by the following formula:

Depreciation Rate = Accelerator *Straight Line Rate

Straight-line Depreciation Rate = 1/4 = 0.25 = 25%
Declining Balance Rate = 2*25% = 50%

Diminishing balance method – Cost of equipment*Depreciation rate

Year 2014= ($152000*50%)*9 months/12 months

= ($76000*9 months)/12 months

= $57000

Book value at the end of year 2014 = $152000-$57000 = $95000

Year 2015 Depreciation expense = $95000*50%= $47500

Book value at the end of year 2015 = $95000-$47500 =$47500

Year 2016 Depreciation expense = $47500*50%= $23750

Book value at the end of year 2016 = $47500-$23750 =$23750

Year 2017 Depreciation expense = $23750*50%= $11875

Book value at the end of year 2017 = $23750-$11875 =$11875

Year 2018 Depreciation expense = $14000-$11875 =$2125

3)- The depreciation expense for 2014 = $19500.

The depreciation expense for 2015 = $30000.

The depreciation expense for 2016 = $37500.

The depreciation expense for 2017 = $27000.

The depreciation expense for 2018 = $24000.

Explanation- Units of production method:-Annual depreciation expense per unit- Machine B

=Cost – salvage /Total units

=($152000-$14000)/9200 hours

=$15 per hour

Depreciation expense in year 2014= Depreciation expense per hour* Hours used

                                              =$15 per hour *1300 hours

                                                 =$19500

Depreciation expense in year 2015= Depreciation expense per hour* Hours used

                                                =$15 per hour *2000 hours

                                                 =$30000

Depreciation expense in year 2016= Depreciation expense per hour* Hours used

                                                =$15 per hour *2500 hours

                                                 =$37500

Depreciation expense in year 2017= Depreciation expense per hour* Hours used

                                                =$15 per hour *1800 hours

                                                 =$27000

Depreciation expense in year 2018= Depreciation expense per hour* Hours used

                                                =$15 per hour *1600 hours

                                                 =$24000